OpenStudy (anonymous):

plzzz. help. whats -0.2(2.12-h)=8.6

OpenStudy (anonymous):

i need to show my work

OpenStudy (lilg132):

open the brackets first -0.2(2.12-h)=8.6 -0.424 + 0.2h = 8.6 now bring the numbers to one side and leave the letts on the other 0.2h = 8.6 + 0.424 0.2h = 9.024 now to get the value for h we have to get it on its own to do that we have to divide it by the number next to it which is 0.2 and if we divide one side we also have to divide the other side so 0.2h/0.2 = 9.024/0.2 h = 45.12 Dont forget to click best response =)

OpenStudy (anonymous):

thank you nd ofcourse I will!